a bag contains six blue pencils, three red pencils, and three pencils. If two penciles are selected at random without replacement, what is the probability that both are red?
Answer by stanbon(75887)

a bag contains six blue pencils, three red pencils, and three pencils. If two penciles are selected at random without replacement, what is the probability that both are red?
--------------
# of ways to succeed: 3C2 = 3
------
# of random pairs: 12C2 = (12/11)/(1*2) = 66
-----
P(2 red pencils) = 3/66 = 1/22
---------------------------------
Cheers,
Stan H.
======================